# Combo: Co-speech holistic 3D human motion generation and efficient customizable adaptation in harmony

### v1

@@ 4.3.2 Analysis on Identity Transfer
Loss w/o. $\mathcal{L}_{F}$,$\mathcal{L}_{B}$ 34.84 30.57 27.19 0.8442 32.46 2910 109.3
$\lambda_{F}$,$\lambda_{B}$ = 1 34.84 1.412 0.593 0.7978 9.963 5.373 6.407
$\lambda_{F}$,$\lambda_{B}$ = 0.5 34.84 1.190 0.579 0.7999 9.995 5.362 6.358
$\lambda_{F}$,$\lambda_{B}$ = 0.1 34.84 1.468 0.598 0.7904 9.880 5.671 6.544
@@
4.3.3 Tuning Efficiency
[…]
@@ 4.4.1 Ablation on DU-Trans
In Tab. VI(a), we conduct the ablation study on DU-Trans from three aspects, i.e., architecture, loss functions, and Bi-Flow variants. The architecture ablation study employs auxiliary losses with optimal weights but does not utilize the Bi-Flow. The loss ablation study fixes the optimal architecture without including the Bi-Flow. Finally, the last study ablates the Bi-Flow on the optimal architecture incorporated with auxiliary losses.
@@
Analysis on Basic Architecture (Tab. VI(a) Architecture part). The insight of our DU-Trans is to use divided encoders to ensure sufficient exhibition for both face and body, followed by a united decoder to implicitly model their interconnections and to directly predict the combined coefficients for overall high coordination. To verify its effectiveness, we first design two variants, the one with only two divided encoders (row $l=8$, $j=0$, where $l$ is the number of encoder layers and $j$ is the number of decoder layers) and the other with only a united decoder (row $l=0$, $j=8$). The results indicate that neither of them yields satisfactory outputs. The underlying reason may be that the two separate encoders do not account for each other’s influence, which is evident from the significant degradation in FMD. Directly employing a single decoder may cause two distinct distributions to converge toward each other, ensuring harmony but greatly sacrificing the uniqueness of each, especially as indicated by the decline in FGD and MSE. We also give a visualization in Fig. 9 that the motion generated by this method has a minimal amplitude. Then, we explore how to preserve individual differences while achieving a harmonious commonality. As shown in row $l=7$, $j=1$, when the encoder has 7 layers and the decoder has 1 layer, our DU-Trans achieves the best performance.
@@
Analysis on Loss Functions (Tab. VI(a) Loss part). The two auxiliary losses for face and body branches are critical to distilling their respective features and aiding in the joint learning of holistic motion. To verify their effectiveness, we present the quantitative results in the Loss part. It is obvious that the absence of auxiliary loss (row w/o. $\lambda_{F},\lambda_{B}$) leads to a significant decrease in most metrics except for BC and DIV. These two are unreliable when there is a noticeable jitter in the motion. This observation can also be discerned in the fourth row in Fig. 9, i.e., exhibiting unreasonable poses and meaningless movements. Furthermore, we explore the optimal hyperparameters for these losses and find that the optimal set is $\lambda_{F}=\lambda_{B}=0.5$.
@@
Analysis on Bi-Flow (Tab. VI(a) Bi-Flow part). The Bi-Flow design is utilized during the divide phase for the face and body to provide global dynamic cues that enhance their respective performances. As shown in the Bi-Flow part, we examine the impact of hyperparameters regarding Bi-Flow, including its position (first 3 rows in this part) and number (the remaining). Concretely, when we apply this module at the shallow layer $l=1$ and the deep layer $l=6$, the performance does not match that at the intermediate layer $l=3$. From row $\lambda_{F}=\lambda_{B}=0.5$ of the Loss part and $l=3$ of the Bi-Flow part, we observe that models incorporating this bidirectional interaction enhance the performance of both the face and body, particularly the face, which shows particularly significant improvement in all metrics. This is consistent with the operation in EMAGE that uses a unidirectional data flow from body to face, yet we find that the reverse flow can also help learning of the body. Besides, this interacted layer also promotes the coordination of the holistic motion, reflecting on the FMD metric. We also provide a qualitative comparison in Fig. 9 to support the above conclusion. Furthermore, building on the foundation of $l=3$, we increase the number of Bi-Flow layers. As indicated in rows $l=3,4$ and $l=2,3,4$, it is observed that while the metrics do not see a significant improvement, the number of trainable parameters increases substantially, i.e., each this layer causing about 5% parameters. Consequently, we determine that placing a single Bi-Flow module at the third layer provides a better balance between tunable parameters and overall performance.
@@
4.4.2 Ablation on X-Adapter
In this section, we ablate the X-Adapter structure under the emotion transfer task from four aspects, i.e., insert location, condition integration method, scale form, and PEFT variants. The results are summarized in Tab. VI(b), where each row only contains one modification over the full version.
Analysis on Insertion Form and Position (Tab. VI(b) Location part). We explore how to insert the added adapter into the original network by comparing the parallel and sequential instances. As shown in rows Serial and Parallel, the parallel form outperforms the sequential one in all metrics. This could be attributed to the parallel adapter receiving the same input as the sub-layers and directly updating the output, which is a more intuitive and natural design. This approach minimally impacts the original model, as the adapter operates independently of the sub-layer outputs. Moreover, we attempt to reduce the use of adapters to reduce the number of trainable parameters. However, from rows Only MHA and Only FFN, we observe that fine-tuning only the adapters inserted into either the FFN or MHA leads to a certain degree of performance decline under the same training settings. Thus, we incorporate our X-Adapter for both the FFN and MHA.
Analysis on Condition (Tab. VI(b) Condition part). It is essential to incorporate the corresponding conditions to control the fine-tuning of emotions or identities. As shown in this part, when the injection of conditions is removed (row w/o. X), all metrics exhibit degradation, indicating that conditions can guide the network to better express the desired information. Moreover, the absence of conditions can also lead to a loss of editing capabilities during the inference phase. We further explore how to integrate the conditions. An intuitive approach is to use a stylization method [71], mapping the conditions into the scale and shift factors to affect the original features, as shown in row Stylization. Although introducing this module enhances performance, it significantly increases the number of trainable parameters, i.e., from 4M to 8.14M. To balance the performance and cost, we continue to experiment with an extreme case, directly adding the condition to the original features without adding any trainable parameters. To our delight, we find that this simple method can achieve comparable results. Therefore, we employ the addition manner to exert the influence of conditions.
Analysis on Scale Form (Tab. VI(b) Scale part). We introduce a dynamic scale mechanism (Dy-Scale) in X-Adapter to dynamically adjust the original features by considering the significance score of the tunable features. To verify its effectiveness, we conduct experiments on two ablated variations: the fixed scalar scale which is set to 1.0 (row Scalar-1.0), and the learnable scalar scale which is initialized by 1.0 (row L-Scalar). As shown in this part, our Dy-Scale yields the best performance with a negligible increase in tunable parameters, from 3.99M to 4.00M. Thus, for such complex temporal motions, learning an adaptive weight for each frame is an intuitive and effective strategy. We further visualize the ratio of updated tokens in each layer for both MHA and FFN to explore the mechanism of Dy-Scale. As shown in Fig. 10, the same audio input exhibits similar changes across different layers of MHA and FFN under various emotional conditions, but the specific ratios are entirely distinct. The neutral output from the fine-tuned model has only a few activated tokens because it shares similar motion patterns as the neutral output from the pre-trained model. In contrast, the emotional output is entirely different, which can also be discerned from the adjustment ratio.
Analysis on Other PEFT Methods (Tab. VI(b) PEFT part). To further prove the effectiveness of our proposed X-Adapter, we compare it with several PEFT approaches, i.e., Prefix Tuning [45] and LoRA [47], under the same training time and data. Specifically, prefix tuning involves prepending 64 learnable tokens in the temporal dimension and adding the conditions to them. From row Prefix Tuning, it is evident that this manner does not achieve effective transfer. While the number of trainable parameters is minimal, it cannot be ignored that an increase in the number of tokens also leads to a substantial increase in memory consumption. Besides, to ensure a fair comparison, we adapt Dy-Scale and the condition in the same manner as LoRA. We set the rank $r$ for both to be 64 thus the tunable parameters are the same. By comparing rows LoRA-r64 and Adapter-r64, the overall performance of the X-Adapter is superior to that of LoRA. Furthermore, we attempt to increase the rank to 128. As observed in rows Adapter-r128 and Adapter-r64, while the number of trainable parameters has increased, there is a noticeable improvement in overall performance. The potential reason may be that complex motion patterns require more parameters to be well-fitted during the fine-tuning phase. Experimentally, we employ the proposed X-Adapter with rank set to 128 in the full version.

### v2

@@ 4.3.2 Analysis on Identity Transfer
[…]
4.3.3 User Study
Similarly, we use the same methodology in Sec. 4.2.3 and conduct the appropriateness tests following the mismatching strategy to verify the effectiveness of X-Adapter on emotion and identity transfer. Our method for this evaluation is the finetuning version. We also include TalkSHOW and EMAGE in the test to demonstrate not only that Combo can generalize to new speakers and emotions, but also that these other models cannot.
For emotion, we sample 2 clips from each emotional speech segment of BEAT-S2, obtaining 14 motion videos (2 segments $\times$ 7 emotion styles). Each is paired with a video generated from a neutral speech segment to form a mismatched pairs. In the emotion generalization test, participants are asked to “Please indicate which motion is generated for the given emotion” (covering 7 emotional styles, excluding neutral).
For identity, we use 4 identities (Speaker-1, -2, -11, -23). Each identity forms 3 mismatched pairs with the remaining identities, such as 1-2, 1-11, 1-23, where the first number represents the given identity in question. This allows us to construct 12 combinations (4 identities $\times$ 3 pairs), with each identity is included 6 times. Concretely, we sample 6 speech segments from each identity, with each of the 12 combinations appearing once. In this test, participants are asked to “Please indicate which motion is generated for the given identity”.
Notably, the videos in both emotion and identity experiments are muted. Besides, the training pages in the experiments include not only the fixed pair videos but also example videos for the 7 emotions and 4 identities that users need to identify.
As a result, 38 and 39 subjects pass the filtering checks, respectively. The results are shown in Tab. III columns 4 and 5. For baselines like EMAGE, the officially released model has relatively weak emotion control, which is realized through the input of emotional speech segments. Its ability to transfer speaker-ID is virtually non-existent due to identity-specific training. The above description can be consistently inferred from their low scores, indicating that their motions across different emotions and identities are difficult to distinguish. Conversely, our results are easily identifiable, demonstrating that our method can generalize to new speakers and emotions, which is attributed to the fast-finetuning property of the proposed X-Adapter.
@@ 4.3.4 Tuning Efficiency
Loss w/o. $\mathcal{L}_{F}$$\mathcal{L}_{B}$ 34.84 30.57 27.19 0.8442 32.46 2910 109.3
$\lambda_{F}$$\lambda_{B}$ = 1 34.84 1.412 0.593 0.7978 9.963 5.373 6.407
$\lambda_{F}$$\lambda_{B}$ = 0.5 34.84 1.190 0.579 0.7999 9.995 5.362 6.358
$\lambda_{F}$$\lambda_{B}$ = 0.1 34.84 1.468 0.598 0.7904 9.880 5.671 6.544
@@
Figure 10: Qualitative ablations for DU-Trans. The full version of DU-Trans ($l=7,j=1,w/\mathcal{L}_{B},\mathcal{L}_{F},w/\text{Bi-Flow}$) produces shows more synchronized and meaningful motions than other ablations.
@@ 4.4.1 Ablation on DU-Trans
In Tab. VI(b)(a), we conduct the ablation study on DU-Trans from three aspects, i.e., architecture, loss functions, and Bi-Flow variants. The architecture ablation study employs auxiliary losses with optimal weights but does not utilize the Bi-Flow. The loss ablation study fixes the optimal architecture without including the Bi-Flow. Finally, the last study ablates the Bi-Flow on the optimal architecture incorporated with auxiliary losses.
@@
Analysis on Basic Architecture (Tab. VI(b)(a) Architecture part). The insight of our DU-Trans is to use divided encoders to ensure sufficient exhibition for both face and body, followed by a united decoder to implicitly model their interconnections and to directly predict the combined coefficients for overall high coordination. To verify its effectiveness, we first design two variants, the one with only two divided encoders (row $l=8$, $j=0$, where $l$ is the number of encoder layers and $j$ is the number of decoder layers) and the other with only a united decoder (row $l=0$, $j=8$). The results indicate that neither of them yields satisfactory outputs. The underlying reason may be that the two separate encoders do not account for each other’s influence, which is evident from the significant degradation in FMD. Directly employing a single decoder may cause two distinct distributions to converge toward each other, ensuring harmony but greatly sacrificing the uniqueness of each, especially as indicated by the decline in FGD and MSE. We also give a visualization in Fig. 10 that the motion generated by this method has a minimal amplitude. Then, we explore how to preserve individual differences while achieving a harmonious commonality. As shown in row $l=7$, $j=1$, when the encoder has 7 layers and the decoder has 1 layer, our DU-Trans achieves the best performance.
@@
Analysis on Loss Functions (Tab. VI(b)(a) Loss part). The two auxiliary losses for face and body branches are critical to distilling their respective features and aiding in the joint learning of holistic motion. To verify their effectiveness, we present the quantitative results in the Loss part. It is obvious that the absence of auxiliary loss (row w/o. $\lambda_{F},\lambda_{B}$) leads to a significant decrease in most metrics except for BC and DIV. These two are unreliable when there is a noticeable jitter in the motion. This observation can also be discerned in the fourth row in Fig. 10, i.e., exhibiting unreasonable poses and meaningless movements. Furthermore, we explore the optimal hyperparameters for these losses and find that the optimal set is $\lambda_{F}=\lambda_{B}=0.5$.
@@
Analysis on Bi-Flow (Tab. VI(b)(a) Bi-Flow part). The Bi-Flow design is utilized during the divide phase for the face and body to provide global dynamic cues that enhance their respective performances. As shown in the Bi-Flow part, we examine the impact of hyperparameters regarding Bi-Flow, including its position (first 3 rows in this part) and number (the remaining). Concretely, when we apply this module at the shallow layer $l=1$ and the deep layer $l=6$, the performance does not match that at the intermediate layer $l=3$. From row $\lambda_{F}=\lambda_{B}=0.5$ of the Loss part and $l=3$ (including two unidirectional variants body-to-face Uni-BF and face-to-body Uni-FB) of the Bi-Flow part, we observe that the unidirectional variants shows enhancement on the target component and overall improvement, while incorporating bidirectional interaction achieves more comprehensive gains. This demonstrates that a bidirectional connection is superior, creating a synergistic effect where the body provides global context to the face (shown in EMAGE) and the face provides fine-grained cues back to the body (shown in DiffSHEG), leading to mutual reinforcement for holistic harmony. We also provide a qualitative comparison in Fig. 10 to support the above conclusion. Furthermore, building on the foundation of $l=3$, we increase the number of Bi-Flow layers. As indicated in rows $l=3,4$ and $l=2,3,4$, it is observed that while the metrics do not see a significant improvement, the number of trainable parameters increases substantially, i.e., each this layer causing about 5% parameters. Consequently, we determine that placing a single Bi-Flow module at the third layer provides a better balance between tunable parameters and overall performance.
4.4.2 Ablation on X-Adapter
In this section, we ablate the X-Adapter structure under the emotion transfer task from four aspects, i.e., insert location, condition integration method, scale form, and PEFT variants. The results are summarized in Tab. VI(b)(b), where each row only contains one modification over the full version.
Analysis on Insertion Form and Position (Tab. VI(b)(b) Location part). We explore how to insert the added adapter into the original network by comparing the parallel and sequential instances. As shown in rows Serial and Parallel, the parallel form outperforms the sequential one in all metrics. This could be attributed to the parallel adapter receiving the same input as the sub-layers and directly updating the output, which is a more intuitive and natural design. This approach minimally impacts the original model, as the adapter operates independently of the sub-layer outputs. Moreover, we attempt to reduce the use of adapters to reduce the number of trainable parameters. However, from rows Only MHA and Only FFN, we observe that fine-tuning only the adapters inserted into either the FFN or MHA leads to a certain degree of performance decline under the same training settings. Thus, we incorporate our X-Adapter for both the FFN and MHA.
Analysis on Condition (Tab. VI(b)(b) Condition part). It is essential to incorporate the corresponding conditions to control the fine-tuning of emotions or identities. As shown in this part, when the injection of conditions is removed (row w/o. X), all metrics exhibit degradation, indicating that conditions can guide the network to better express the desired information. Moreover, the absence of conditions can also lead to a loss of editing capabilities during the inference phase. We further explore how to integrate the conditions. An intuitive approach is to use a stylization method [73], mapping the conditions into the scale and shift factors to affect the original features, as shown in row Stylization. Although introducing this module enhances performance, it significantly increases the number of trainable parameters, i.e., from 4M to 8.14M. To balance the performance and cost, we continue to experiment with an extreme case, directly adding the condition to the original features without adding any trainable parameters. To our delight, we find that this simple method can achieve comparable results. Therefore, we employ the addition manner to exert the influence of conditions.
Analysis on Scale Form (Tab. VI(b)(b) Scale part). We introduce a dynamic scale mechanism (Dy-Scale) in X-Adapter to dynamically adjust the original features by considering the significance score of the tunable features. To verify its effectiveness, we conduct experiments on two ablated variations: the fixed scalar scale which is set to 1.0 (row Scalar-1.0), and the learnable scalar scale which is initialized by 1.0 (row L-Scalar). As shown in this part, our Dy-Scale yields the best performance with a negligible increase in tunable parameters, from 3.99M to 4.00M. Thus, for such complex temporal motions, learning an adaptive weight for each frame is an intuitive and effective strategy. We further visualize the ratio of updated tokens in each layer for both MHA and FFN to explore the mechanism of Dy-Scale. As shown in Fig. 11, the same audio input exhibits similar changes across different layers of MHA and FFN under various emotional conditions, but the specific ratios are entirely distinct. The neutral output from the fine-tuned model has only a few activated tokens because it shares similar motion patterns as the neutral output from the pre-trained model. In contrast, the emotional output is entirely different, which can also be discerned from the adjustment ratio.
Analysis on Other PEFT Methods (Tab. VI(b)(b) PEFT part). To further prove the effectiveness of our proposed X-Adapter, we compare it with several PEFT approaches, i.e., Prefix Tuning [47] and LoRA [hu2021lora], under the same training time and data. Specifically, prefix tuning involves prepending 64 learnable tokens in the temporal dimension and adding the conditions to them. From row Prefix Tuning, it is evident that this manner does not achieve effective transfer. While the number of trainable parameters is minimal, it cannot be ignored that an increase in the number of tokens also leads to a substantial increase in memory consumption. Besides, to ensure a fair comparison, we adapt Dy-Scale and the condition in the same manner as LoRA. We set the rank $r$ for both to be 64 thus the tunable parameters are the same. By comparing rows LoRA-r64 and Adapter-r64, the overall performance of the X-Adapter is superior to that of LoRA. Furthermore, we attempt to increase the rank to 128. As observed in rows Adapter-r128 and Adapter-r64, while the number of trainable parameters has increased, there is a noticeable improvement in overall performance. The potential reason may be that complex motion patterns require more parameters to be well-fitted during the fine-tuning phase. Experimentally, we employ the proposed X-Adapter with rank set to 128 in the full version.
@@ 4.4.2 Ablation on X-Adapter
Table VII: Average scores of user study with 95% confidence intervals. The left part is for DU-Trans and the right is for X-Adapter.
Human-likeness Synchronized Coordinated Generalized-E
[…]
4.4.3 User Study
We further conduct the user study of ablation studies as additional validation for the aforementioned objective conclusions. Similarly, we use the same approach as in Sec. 4.2.3.
For ablation on DU-Trans, we select four variants, Exp 1 corresponds to row $l=0,j=8$, Exp 2 to row $l=8,j=0$, Exp 3 to row $l=7,j=1$, and Exp 4 to row w/o. $\lambda_{F},\lambda_{B}$, respectively, while Ours corresponds to row $l=3$ in Tab. VI(b)(a). We adopt the human-likeness (20 combinations appear once, 20 video pairs, 40 qualified participants), synchronized (8 video pairs for each method, 37 qualified participants), and coordinated (8 video pairs for each method, 38 qualified participants) three user studies. The average score shown in the left part of Tab. VII. Notably, we do not include the mismatching tests for Exp 4 since it fails to produce plausible human motions, as evidenced in Fig. 10 row 4. Specifically, comparing Exps 1–3 shows that our divided encoder and unified decoder preserve individual differences while achieving commonality. Moreover, comparing Exps 3 and 4 reveals that the auxiliary losses for face and body help capture their unique features. Lastly, comparing Exp 3 and Ours confirms that the Bi-Flow mechanism further boosts performance.
For ablation on X-Adapter, to align with the Tab. VI(b), we only evaluate emotion transfer capability through the Generalized-E test (8 video pairs for each method, 38 qualified participants). The average score shown in the right part of Tab. VII. Exp 5 corresponds to row Serial, Exp 6 to row w/o. X, Exp 7 to row L-Scalar, and Exp 8 to row Adapter-r64, while Ours corresponds to row Adapter-r128 in Tab. VI(b)(b). Our method’s mismatched percentage is higher than that of the four variants. This indicates that: 1) The parallel insertion method outperforms the serial one. 2) The absence of an emotional condition significantly degrades transfer accuracy. 3) Our proposed dynamic learning with an adaptive weight per frame (Dy-Scale) is superior to using a global weight (L-Scalar). 4) Increasing the adapter’s rank enhances performance, respectively.
@@ 5 Conclusion
Acknowledgements
This work is supported in part by National Science and Technology Major Project under contract No. 2021ZD0109902, in part by the National Natural Science Foundation of China under contract No. 62171256.